If your car is traveling 70 mi/he and your cars direction changes 2°, how far to the right would your car move
1
Expert's answer
2020-11-23T10:27:05-0500
"v_y=v\sin{2}\\v_y=70\sin{2}=2.4\frac{mi}{h}"